. Average Rate of Return—New Product Pocket Pilot Inc. is considering an investment in new equipment...

.

Average Rate of Return—New Product

Pocket Pilot Inc. is considering an investment in new equipment that will be used to manufacture a mobile communications device. The device is expected to generate additional annual sales of 4,100 units at $319 per unit. The equipment has a cost of $457,600, residual value of $34,400, and an eight-year life. The equipment can only be used to manufacture the device. The cost to manufacture the device is shown below.

Cost per unit:
Direct labor $54
Direct materials 209
Factory overhead (including depreciation) 35.6
Total cost per unit $298.6

Determine the average rate of return on the equipment. If required, round to the nearest whole percent.
Answer #1

ARR = (Average Profit per year / Average investment) × 100

Average profit per year = (Sale price – Cost per unit) × Number of units per year

                                    = (319 – 298.60) × 4,100

                                    = $20.40 × 4,100

                                    = $83,640

Average investment = (Cost + Residual value) / 2

                                    = (457,600 + 34,400) / 2

                                    = 492,000 / 2

                                    = $246,000

ARR = (83,640 / 246,000) × 100

            = 34% (Answer)